# Nonflat bands and chiral symmetry in magic-angle twisted bilayer graphene

### v1

@@ Abstract
In this work, we derive effective theories for magic-angle twisted bilayer graphene (MATBG) based on the mean-field normal state at charge neutrality of a tight-binding model, thereby including the renormalizations coming from integrating out high-energy modes. In these models, the flat bands display a sizable increase of the bandwidth, and the corresponding wavefunctions flow towards the limit of perfect particle-hole symmetry and sublattice polarization. We further represent the flat bands in the ’vortex Chern’ basis and also discuss the implications on the dynamics regarding the ’flat’ and ’chiral’ symmetries of MATBG, as manifested in the symmetry-broken states at neutrality.
@@
Introduction.
Magic-angle twisted bilayer graphene (MATBG), as a type of moiré material [1, 2], hosts a mesoscopic unit cell and flat electronic bands at charge neutrality. It represents a tunable platform for the study of strong interactions and topology [3, 4, 5], as well as their interplay, thus attracting great attention in the recent years. Experimental and theoretical efforts have unveiled a rich phase diagram, exhibiting in a single material phenomena like correlated states [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27], Kondo physics [28, 29, 30, 31, 32, 33, 34] and superconductivity [35, 36, 17, 18, 19, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46], among others [47, 48, 49, 50, 51].
The large number of atoms in the moiré unit cell, and consequently large number of electronic bands, calls for the construction of effective theories of MATBG with a small number of degrees of freedom, that ought to capture accurately the dispersion, symmetry and topology of the low-energy bands. Among such theories there are continuum models [52, 53, 54, 55], including the seminal Bistritzer-MacDonald (BM) model [56], Wannier orbital [57, 58, 59, 60, 61, 62, 63, 64, 65], momentum space [66, 67] and the topological heavy fermion (THF) model [68].
Their very small bandwidth renders the flat bands prone to distortions coming from otherwise minor perturbations, like lattice relaxation [54, 53, 69, 70] or strain [69, 71, 72]. In this respect, an often overlooked piece of the physics is the interacting effect of the high-energy degrees of freedom removed in the effective theory, but nonetheless implicit in the ground state. We argue that virtual processes between the ’remote’ (high-energy) and ’active’ (low-energy) modes are significant and must be accounted for [73], at least approximately. An earlier work addressed this problem using a Renormalization Group approach [74].
In this Letter, we construct a normal state at charge neutrality that preserves the symmetries and incorporates the Hartree-Fock potentials of the occupied graphene $\pi$ orbitals self-consistently; any low-energy theory of MATBG ought to reproduce the dispersion and wavefunctions of the normal, rather than the non-interacting, state. In this work we obtained effective models for MATBG using a many-body projection scheme onto the central bands of the normal state, sketched in Fig. 1.
In the renormalized system the dispersion widens drastically, with the flat bandwidth reaching several times the non-interacting one and becoming comparable to the energy scale of the Coulomb interactions among flat electrons. Furthermore, we find that the flat-band Hilbert space flows towards the limit of perfect particle-hole symmetry and sublattice polarization.
At the same time, there are important modifications of the symmetry of MATBG [6, 75, 68]. The $U(4)_{\text{flat}}$ symmetry of the ’nonchiral-flat’ limit, that relies on the existence of very flat bands, is strongly broken; instead, the ’chiral-nonflat’ $U(4)_{\text{chiral}}$ group prevails as an approximate symmetry independently of the bandwidth. Such properties become apparent in the symmetry-broken phases. Lastly, we comment on the experimental imprints of our results on the recently observed spiral states [76].
@@ Introduction.
Figure 2: (a) Band structure of the symmetric state for different interaction strengths, compared to the non-interacting (bare) bands. (b) Hilbert space of the bare and renormalized ($\epsilon_{r}=10,$ $U=4$ eV) flat bands. In the top row we show the positive eigenvalue of the sublattice operator, $\overline{\sigma}_{z}$, (sublattice polarization). The remaining eigenvalues are obtained from $180^{\circ}$ rotations and sign changes. In the bottom rows we plot the two distinct singular values of the projected particle-hole operator $C_{2z}\overline{P}$, which are degenerate in pairs.
Symmetric state and many-body projection.
@@
We start with a tight-binding model for MATBG with one orbital per carbon atom. The twist angle is chosen to $\theta=1.05^{\circ}$, resulting in a unit cell with 11908 atoms, and in-plane atomic relaxations are included following Ref. [77]. The crystalline symmetries of the lattice include six-fold rotations around the $z$ axis, $C_{6z}$, and two-fold rotations around the $y$ axis between the layers, $C_{2y}$. Below we will be addressing the products $C_{3z}=C_{6z}^{2}$, $C_{2z}=C_{6z}^{3}$ and $C_{2x}=C_{2y}C_{2z}$ also. Spinless time-reversal, $\mathcal{T}$, acting as complex conjugation, is a symmetry if we neglect spin-orbit coupling.
@@
We find a self-consistent solution of ${H}$ in mean-field (Hartree-Fock) theory that preserves spin rotations, time-reversal symmetry and the crystallographic symmetries (see the Supplementary Material 111see Supplementary Material, which includes the references [94, 77, 95, 80, 81, 90, 96, 75, 6, 85, 4, 5, 84, 54, 69, 87, 88, 68, 97, 92, 15, 41, 98, 99, 100, 101, 102, 103] for details). In Fig. 2(a) we plot the band structures of the symmetric state with $\epsilon_{r}=10$, $U=4$ eV and $\epsilon_{r}=50$, $U=0.5$ eV (the weak coupling limit corresponds to the physical situation if the self-screening is strong [83]), and compare them with the bare band structure. The flat bandwidth increases approximately linearly with $\epsilon_{r}^{-1}$, growing from $8.7$ meV to $15.3$ meV to $48.4$ meV in the bare, $\epsilon_{r}=50$ and $\epsilon_{r}=10$ bands, respectively. Another relevant quantity is the gap between the flat and remote bands, controlling the validity of the flat-band projection [66]. The gaps to the remote bands increase from $20.2$ to $26.8$ meV on the electron side and from $23.9$ to $30.8$ meV on the hole side in the bare and $\epsilon_{r}=10$ bands, respectively.
Not only the energies, but also the wave functions are renormalized. We study the properties of Hilbert space of the flat bands in Fig. 2(b). We plot the sublattice polarization, this is, the eigenvalues of the sublattice operator projected to the flat bands, $\overline{\sigma}_{z}$, and the singular values of the ’particle-hole’ operator [84, 5, 54, 69] projected to the flat bands, $C_{2z}\overline{P}$. The system flows towards a perfectly sublattice polarized and particle-hole symmetric limit with eigenvalues (singular values) equal to $1$.
Using the projector onto many-body excitations of the lowest $n_{B}$ bands of the normal state, $\mathcal{P}$, we obtain an effective model for MATBG with Hamitonian ${\mathcal{H}}=\mathcal{P}H\mathcal{P}=\mathcal{H}_{0}+{\mathcal{H}}_{\text{int}}$,
@@
In this expression, the operator $c^{\dagger}_{\bm{k}ns}(c_{\bm{k}ns})$ creates (annihilates) an electron with momentum $\bm{k}$, band $n$ and spin $s$ in the mean-field basis, and $n,m$ are restricted to the central $n_{B}$ bands, $n,m=\{-n_{B}/2,...,n_{B}/2\}$. In both cases, the $n_{B}$ lowest bands are gapped from the remaining bands at all $\bm{k}$, so the projector is well-defined.
The matrix ${h}(\bm{k})_{nm}$ contains the dispersion of the normal state plus a counter-term to avoid double-counting the valence bands $20<n\leq-n_{B}/$ 2, and a term related to the order of the fermionic operators, see the Supplementary Materials [82] for details. On the other hand, ${\rho}_{\bm{r}s}$ is the density relative to charge neutrality,
@@
where $\langle\bm{r}|\bm{k}n\rangle$ is the wave function amplitude of mode $c^{\dagger}_{\bm{k}ns}$ at position $\bm{r}$ and again $n,m=\{-n_{B}/2,...,n_{B}/2\}$. The choice of $\rho_{0\bm{r}}$ is consistent with the approximate particle-hole symmetry.
The band structure of $\mathcal{H}_{0}$ is shown in Fig. 3(a) with $\epsilon_{r}=10$, $U=4$ eV, for $n_{B}=4$ and $20$. In both cases, the dispersion is only slightly narrower than in the normal state.
Vortex Chern gauge and nonflat bands.
@@
The ’Chern gauge’ [6, 75] fixes at each $\bm{k}$ the relative phase between bands with the same Chern number, leaving an ambiguity of the form $\text{exp}(i\phi(\bm{k})\sigma_{z}\tau_{z})$ [6, 86, 13] (see the Supplementary Materials [82] for details on the gauge fixing and symmetry representations). We fix $\phi(\bm{k})$ by requiring that the Wannier functions obtained from the Chern bands are maximally localized [87, 88], following the algorithm of Ref. [87]; the Bloch states are periodic and smooth except at the $\Gamma_{M}$ point, where the Berry connection acquires a vortex with a winding number equal to the Chern number. This ’vortex Chern gauge’ is defined up to a global sign of all wave functions.
@@
Symmetry and correlated states.
Here, we comment on the consequences of our renormalization procedure on the symmetry in the $n_{B}=4$ system. For a detailed analysis of the symmetries of MATBG, including larger $n_{B}$, we refer the reader to the Supplementary Materials [82] as well as to Refs. [75, 6, 68].
With respect to the interaction, the breaking of $\mathcal{G}_{\text{chiral}}$ and $\mathcal{G}_{\text{flat}}$ can be assessed by the departure from one of the eigenvalues and singular values of $\overline{\sigma}_{z}$ and $C_{2z}\overline{P}$, respectively, displayed in Fig. 2(b). They increase in the symmetric state, thus enhancing the symmetry of ${\mathcal{H}}_{\text{int}}$; this phenomenon was already noted in Ref. [74] for $\overline{\sigma}_{z}$.
Invariably in all phases, the kinetic energy penalty induces $Q(\bm{k})\sim\sigma_{x},\sigma_{y}\tau_{z}$ around the $\Gamma_{M}$ point, coming from the lower bands of $\mathcal{H}_{0}$. This is apparent in Fig. 4(b) where we plot the $Q(\bm{k})$ matrix of KIVC and OP. Concurrently, the wide dispersion enforces TIVC, VP and SP to be gapless, while KIVC, OP and QAH are insulating, see Fig. 4(a). These (semi)metallic states are destabilized by $\mathcal{H}_{0}$ as compared to the others. Moreover, the approximate chiral symmetry, $|\text{KIVC}\rangle\approx\text{exp}(i\pi/4\ \mathcal{G}_{\text{chiral}})|\text{OP}\rangle$, $|\text{VP}\rangle\approx\text{exp}(i\pi/4\ \mathcal{G}_{\text{chiral}})|\text{TIVC}\rangle$ is manifest in Fig. 4(a) and (b).
@@
Figure 4: (a) Band structure of the Kramers intervalley coherent (KIVC), orbital polarized (OP), time-reversal invariant intervalley coherent (TIVC) and valley polarized (VP) states. The approximate chiral symmetry relating KIVC and OP, and TIVC and VP is manifest. (b) Order parameters, defined by $Q(\bm{k})=\sum_{ij}\langle\sigma_{i}\tau_{j}(\bm{k})\rangle\sigma_{i}\tau_{j}$, of the KIVC and OP states. The quantities $\langle\sigma_{y}\tau_{y}(\bm{k})\rangle$ and $\langle\sigma_{z}(\bm{k})\rangle$ are related by a chiral rotation, and the quantities $\langle\sigma_{x}(\bm{k})\rangle$ and $\langle\sigma_{y}\tau_{z}(\bm{k})\rangle$ are trivial under chiral rotations corresponding to filling the valence bands of $\mathcal{H}_{0}$. The coupling strength is $\epsilon_{r}=10$, $U=4$ eV.
Discussion.
In this work, we performed large scale (11908 atoms per unit cell) computations in an atomistic model of MATBG and obtained a mean-field symmetric state at charge neutrality. In this state, the dispersion and the wave functions are renormalized such that the electrons are decoupled at the one-loop (Hartree-Fock) level. We argue that effective models with a reduced bandwidth reproducing this symmetric phase take account of the many-body effects coming from the removed degrees of freedom, as opposed to models obtained from a naive truncation of the bare Hilbert space.
We have constructed effective theories using a many-body projection onto the lowest $4$ and $20$ bands per spin of the symmetric state. The $4$-band model corresponds to the flat band projection, and the $20$-band model describes an energy window of about ±$\pm 250$ meV. In these theories the flat bands widen drastically, with a bandwidth comparable to the interaction scale, demonstrating the relevance of the renormalizations induced by integrating out the remote bands
The widening of the flat bands is directly related to the breaking of the $U(4)_{\text{flat}}$ symmetry. Instead, we find that the chiral group $U(4)_{\text{chiral}}$ is an approximate symmetry of the system, broken at a scale of some tens of meV$/\epsilon_{r}$.

### v3

@@ Abstract
In this work, we study an interacting tight-binding model of magic-angle twisted bilayer graphene (MATBG), with a twist angle of $1.05^{\circ}$. We derive effective theories based on a mean-field normal state at charge neutrality, thereby including the renormalizations coming from integrating out high-energy modes. In these theories, the flat bands display a sizable increase of the bandwidth, suggesting the renormalization of the magic angle. Additionally, the corresponding wavefunctions flow towards the limit of perfect particle-hole symmetry and sublattice polarization (the ’chiral’ limit). We further represent the flat bands in the ’vortex Chern’ basis and discuss the implications on the dynamics, regarding the ’flat’ and ’chiral’ symmetries of MATBG, as manifested in the symmetry-broken states at neutrality.
@@
I Introduction
@@
Magic-angle twisted bilayer graphene (MATBG), as a type of moiré material [1, 2], hosts a mesoscopic unit cell and flat electronic bands at charge neutrality. It represents a tunable platform for the study of strong interactions and topology [3, 4, 5], as well as their interplay, thus attracting great attention in the recent years. Experimental and theoretical efforts have unveiled a rich phase diagram, exhibiting in a single material phenomena like correlated states [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28], Kondo physics [29, 30, 31, 32, 33, 34, 35] and superconductivity [36, 37, 17, 18, 19, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47], among others [48, 49, 50, 51, 52, 53].
The large number of atoms in the moiré unit cell, and consequently large number of electronic bands, calls for the construction of effective theories of MATBG with a small number of degrees of freedom, that ought to capture accurately the dispersion, symmetry and topology of the low-energy bands. Among such theories there are continuum models [54, 55, 56, 57], including the seminal Bistritzer-MacDonald (BM) model [58], Wannier orbital [59, 60, 61, 62, 63, 64, 65, 66, 67], momentum space [68, 69, 44] and the topological heavy fermion (THF) model [70].
Their very small bandwidth renders the flat bands prone to distortions coming from otherwise minor perturbations, like lattice relaxation [56, 55, 71, 72] or strain [71, 73, 74]. In this respect, an often overlooked piece of the physics is the interacting effects of the high-energy degrees of freedom that are removed in the effective theory, but remain implicit in the ground state. We argue that virtual processes between the ’remote’ (high-energy) and ’active’ (low-energy) modes are significant and must be accounted for [75], at least approximately. An earlier work addressed this problem using a Renormalization Group approach [76].
In this Article, we consider twisted bilayer graphene with a twist of $1.05^{\circ}$, around the nominal magic angle [56]. We obtain a normal state at charge neutrality that preserves the symmetries and incorporates the Hartree-Fock potentials of the occupied graphene $\pi$ orbitals self-consistently; any low-energy theory of MATBG ought to reproduce the dispersion and wavefunctions of the normal, rather than the non-interacting, state. We construct effective models for MATBG using a many-body projection scheme onto the central bands of this symmetric state, sketched in Fig. 1.
In the renormalized system the dispersion widens drastically with the flat bandwidth reaching several times the non-interacting one, becoming comparable to the energy scale of the Coulomb interactions among flat electrons. This suggests the tuning of the magic angle (or magic ’range’ [54]) by the interactions, in a similar way to lattice relaxation [77, 54]. Furthermore, the flat-band Hilbert space flows towards the limit of perfect particle-hole symmetry and sublattice polarization (the chiral limit [3, 78, 76]).
At the same time, there are important modifications of the symmetry of MATBG [6, 79, 70]. The $U(4)_{\text{flat}}$ symmetry of the ’nonchiral-flat’ limit, that relies on the existence of very flat bands, is strongly broken; instead, the ’chiral-nonflat’ $U(4)_{\text{chiral}}$ group prevails as an approximate symmetry independently of the bandwidth. Such properties become apparent in the symmetry-broken phases at charge neutrality.
@@
Figure 2: (a) Band structure of the symmetric state for different interaction strengths, compared to the non-interacting (bare) bands. The inset shows the Brillouin zone and the momentum lines plotted. (b) Hilbert space of the bare and renormalized ($\epsilon_{r}=10,$ $U=4$ eV) flat bands. In the top row we show the positive eigenvalue of the sublattice operator, $\overline{\sigma}_{z}$, (sublattice polarization) as a function of momentum. The remaining eigenvalues are obtained from $180^{\circ}$ rotations and sign changes. In the bottom rows we plot the two distinct singular values of the projected particle-hole operator $C_{2z}\overline{P}$, which are degenerate in pairs.
II Symmetric state and many-body projection
@@
We start with a tight-binding model for MATBG with one orbital per carbon atom; there are 11908 atoms in the unit cell of the $1.05^{\circ}$ twisted structure. In-plane atomic relaxations are included following Ref. [77]. The crystalline symmetries of the lattice include six-fold rotations around the $z$ axis, $C_{6z}$, and two-fold rotations around the $y$ axis between the layers, $C_{2y}$; below we will be addressing the products $C_{3z}=C_{6z}^{2}$, $C_{2z}=C_{6z}^{3}$ and $C_{2x}=C_{2y}C_{2z}$ also. Spinless time-reversal, $\mathcal{T}$, acting as complex conjugation, is a symmetry if we neglect spin-orbit coupling.
@@
We find a self-consistent solution of ${H}$ in mean-field (Hartree-Fock) theory that preserves spin rotations, time-reversal symmetry and the crystallographic symmetries (see Appendix B for details). In Fig. 2(a) we plot the band structures of the symmetric state with $\epsilon_{r}=10$, $U=4$ eV and $\epsilon_{r}=50$, $U=0.5$ eV (the weak coupling limit corresponds to the physical situation if the self-screening is strong [84]), and compare them with the bare band structure. The flat bandwidth increases approximately linearly with $\epsilon_{r}^{-1}$, growing from $8.7$ meV to $15.3$ meV to $48.4$ meV in the bare, $\epsilon_{r}=50$ and $\epsilon_{r}=10$ bands, respectively. Another relevant quantity is the gap between the flat and remote bands, controlling the validity of the flat-band projection [68]; the gaps increase from $20.2$ to $26.8$ meV on the electron side and from $23.9$ to $30.8$ meV on the hole side in the bare and $\epsilon_{r}=10$ bands, respectively.
Not only the energies, but also the wave functions are renormalized. In Fig. 2(b) we study the properties of Hilbert space of the flat bands; we plot the sublattice polarization, this is, the eigenvalues of the sublattice operator projected to the flat bands, $\overline{\sigma}_{z}$, and the singular values of the ’particle-hole’ operator [85, 5, 56, 71] projected to the flat bands, $C_{2z}\overline{P}$. The system flows towards a perfectly sublattice polarized and particle-hole symmetric limit with eigenvalues, or singular values, equal to $1$. The increase of the $\overline{\sigma}_{z}$ eigenvalues, i.e. the flow towards the chiral limit [3, 78], was already noted in Ref. [76].
Using the projector onto many-body excitations of the lowest $n_{B}$ bands (per spin) of the normal state, $\mathcal{P}$, we can obtain an effective model for MATBG with the Hamitonian ${\mathcal{H}}=\mathcal{P}H\mathcal{P}=\mathcal{H}_{0}+{\mathcal{H}}_{\text{int}}$,
@@
In this expression, the operator $c^{\dagger}_{\bm{k}ns}(c_{\bm{k}ns})$ creates (annihilates) an electron on the $n^{\text{th}}$ mean-field band with momentum $\bm{k}$ in the first Brillouin zone and spin $s$, and $n,m$ are restricted to the central $n_{B}$ bands, $n,m=\{-n_{B}/2,...,n_{B}/2\}$. We take $n_{B}=4$ and $20$, for which the band projector is well-defined, see Appendix B for details.
The matrix ${h}(\bm{k})_{nm}$ contains the dispersion of the normal state plus a counter-term to avoid double-counting the valence bands $20<n\leq-n_{B}/$ 2, and a term related to the order of the fermionic operators, see Appendix B for details. On the other hand, ${\rho}_{\bm{r}s}$ is the density relative to charge neutrality,
@@
where $\langle\bm{r}|\bm{k}n\rangle$ is the wave function amplitude of mode $c^{\dagger}_{\bm{k}ns}$ at position $\bm{r}$; again $n,m=\{-n_{B}/2,...,n_{B}/2\}$. The choice of $\rho_{0\bm{r}}$ is consistent with the approximate particle-hole symmetry.
The band structure of $\mathcal{H}_{0}$ is shown in Fig. 3(a) with $\epsilon_{r}=10$, $U=4$ eV, for $n_{B}=4$ and $20$. In both cases, the dispersion is only slightly narrower than in the normal state, evidencing the crucial effects coming from the removed bands in the projected theories.
III Vortex Chern gauge and nonflat bands
@@
The ’Chern gauge’ [6, 79] fixes at each $\bm{k}$ the relative phase between bands with the same Chern number, leaving an ambiguity of the form $\text{exp}(i\phi(\bm{k})\sigma_{z}\tau_{z})$ [6, 87, 13] (see Appendix C for details on the gauge fixing and symmetry representations). We fix $\phi(\bm{k})$ by requiring that the Wannier functions obtained from the Chern bands are maximally localized [88, 89], following the algorithm of Ref. [88]; the Bloch states are periodic and smooth except at the $\Gamma_{M}$ point, where the Berry connection acquires a vortex with a winding number equal to the Chern number. This ’vortex Chern gauge’ is defined up to a global sign of all wave functions.
@@
IV Symmetry and correlated states
Here, we comment on the consequences of our renormalization procedure on the symmetry in the $n_{B}=4$ system. For a detailed analysis of the symmetries of MATBG, including larger $n_{B}$, we refer the reader to Appendix D as well as to Refs. [79, 6, 70].
With respect to the interaction, the breaking of $\mathcal{G}_{\text{chiral}}$ and $\mathcal{G}_{\text{flat}}$ can be assessed by the eigenvalues and singular values of $\overline{\sigma}_{z}$ and $C_{2z}\overline{P}$, respectively, displayed in Fig. 2(b); values less than one reveal the non-symmetry of the Hilbert space. These quantities increase in the renormalized state, thus enhancing the symmetry of ${\mathcal{H}}_{\text{int}}$.
Invariably in all phases, the kinetic energy penalty induces $Q(\bm{k})\sim\sigma_{x},\sigma_{y}\tau_{z}$ around the $\Gamma_{M}$ point, coming from the lower bands of $\mathcal{H}_{0}$. This is apparent in Fig. 4(b) where we plot the $Q(\bm{k})$ matrix of KIVC and OP. Concurrently, the wide dispersion enforces TIVC, VP and SP to be gapless, while KIVC, OP and QAH are insulating, see Fig. 4(a). These (semi)metallic states are destabilized by $\mathcal{H}_{0}$ as compared to the insulators and the NSM. Moreover, the approximate chiral symmetry, $|\text{KIVC}\rangle\approx\text{exp}(i\pi/4\ \mathcal{G}_{\text{chiral}})|\text{OP}\rangle$, $|\text{VP}\rangle\approx\text{exp}(i\pi/4\ \mathcal{G}_{\text{chiral}})|\text{TIVC}\rangle$, is manifest in Fig. 4(a) and (b).
Figure 4: (a) Band structure of the Kramers intervalley coherent (KIVC), orbital polarized (OP), time-reversal invariant intervalley coherent (TIVC) and valley polarized (VP) states. The approximate chiral symmetry relating KIVC and OP, and TIVC and VP is manifest. (b) Order parameters, defined by $Q(\bm{k})=\sum_{ij}\langle\sigma_{i}\tau_{j}(\bm{k})\rangle\sigma_{i}\tau_{j}$, of the KIVC and OP states. The quantities $\langle\sigma_{y}\tau_{y}(\bm{k})\rangle$ and $\langle\sigma_{z}(\bm{k})\rangle$ are related by a chiral rotation, and the quantities $\langle\sigma_{x}(\bm{k})\rangle$ and $\langle\sigma_{y}\tau_{z}(\bm{k})\rangle$ are trivial under chiral rotations and correspond to filling the valence bands of $\mathcal{H}_{0}$. The coupling strength is $\epsilon_{r}=10$, $U=4$ eV.
V Discussion
In this work, we performed large-scale (11908 atoms per unit cell) computations in an atomistic model of MATBG and obtained a mean-field symmetric state at charge neutrality. In this state, the dispersion and the wave functions are renormalized such that the electrons are decoupled at the one-loop (Hartree-Fock) level. We argue that effective models with a reduced bandwidth reproducing this symmetric phase take account of the many-body effects coming from the removed degrees of freedom, as opposed to models obtained from a naive truncation of the bare Hilbert space.
We have constructed effective theories using a many-body projection onto the lowest $4$ and $20$ bands per spin of the symmetric state. The $4$-band model corresponds to the flat band projection, and the $20$-band model describes an energy window of about ±$\pm 250$ meV.
In these theories the flat bands widen drastically, with a bandwidth comparable to the interaction scale, demonstrating the relevance of the renormalizations induced by integrating out the remote bands. Furthermore, the sublattice polarization of the flat-band subspace increases, showing that the renormalized system approaches the chiral limit of MATBG [3, 78, 76].
The nonflat bands are directly linked to the breaking of the $U(4)_{\text{flat}}$ symmetry. Instead, $U(4)_{\text{chiral}}$ prevails as an approximate symmetry of the system, broken at a scale of some tens of meV$/\epsilon_{r}$, as manifested in the correlated states at charge neutrality.